OpenStudy (anonymous):

Solve 3x2 = –30x + 24

OpenStudy (anonymous):

divide each term by 3 x^2 = -10x + 8 add 10x and -8 to both sides x^2 + 10x - 8 = 0 Now we need to use the quadratic formula

OpenStudy (anonymous):

whats the answer???

OpenStudy (anonymous):

\[\Large \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]

OpenStudy (anonymous):

If a quadratic is of the form \(a{x^2} + bx + c = 0\) then you can plug the appropriate numbers into the quadratic formula (above) to solve.

OpenStudy (anonymous):

but idk how :(

OpenStudy (anonymous):

a = 1 b = 10 c = -8

OpenStudy (anonymous):

i still dont know the answer

OpenStudy (anonymous):

you will need to do some math, adding and multiplying